# 3D Motion Generation for Motion Language Models

### v1

@@ 4.1 Datasets, Metrics and Baselines
We conduct extensive experiments on two widely used large-scale datasets following the existing works plan_posture . The first is the IDEA400 dataset, a high-quality whole-body motion dataset composed of 12.5K clips and 2.6M frames in MotionX motionx , which is utilized to assess OOD performance. The other dataset, HumanML3D humanml3d , comprising 14,616 motion clips and 44,970 text descriptions, is utilized to evaluate in-domain performance. For evaluation metrics, we adopt several widely recognized measures: Frechet Inception Distance (FID), R-recall, MultiModal Distance (MM Dist), Diversity. More details can be found in Appendix A due to the space limit.
@@ 4.3 Main Results
Tables 1 and 2 present a quantitative comparison between VimoRAG and SoTA techniques. For a fair comparison, each experiment is conducted ten times, and we report the results with a 95% confidence interval. As illustrated in Table 1, VimoRAG achieves the best FID score, indicating its strong generalization capability in generating high-fidelity motions in OOD scenarios. According to Table 2, VimoRAG outperforms MotionGPT motiongpt_aaai by a large margin across all metrics when using the same backbone. It achieves the best FID score and competitive performance on other metrics among motion LLMs based on the Phi-3 3.8B backbone. As shown in Figure 7, performance improves steadily with larger retrieval sets, highlighting VimoRAG’s potential to enhance motion LLMs.
@@ 4.4 Ablation Study
Impact of Motion Video Retriever.
To study the the influence of the video retriever, we replace the Gemini-MVR with two other retrievers: one being the fine-tuned InternVideo internvideo model, and the other being random retrieval from the HcVD. The results in Table LABEL:tab:ablation show that random retrieval leads to a notable increase in the FID score, indicating the importance of accurate video priors in generating high-fidelity results. Moreover, replacing the retriever with InternVideo also leads to a rise in FID score, further confirming the effectiveness of Gemini-MVR. It is important to note that the performance of the retrieval model in RAG is influenced by the generation module, hence we also conduct text-to-video retrieval experiments. As shown in Table LABEL:tab:text-to-video, Gemini-MVR achieves an 8.8% increase in R@1 for the human-centric video set (pool size is 1990), and a 16.6% increase in R@1 for the single human-centric video set. These results further validate the effectiveness of Gemini-MVR.
Impact of McDPO.
@@ 4.5 In-depth Discussion
5 Conclusion
@@ A.1 Experimental Settings
MoMask momask 0.048±.004 0.519±.005 0.715±.005 0.809±.004 2.955±.011 9.632±.094
T2M-GPT t2m-gpt 0.112±.006 0.489±.006 0.679±.006 0.774±.004 3.125±.015 9.691±.062
MDM mdm 0.454±.012 0.419±.004 0.606±.004 0.712±.004 3.636±.015 9.449±.136
MotionDiffuse motion_diffuse 0.672±.025 0.492±.004 0.685±.003 0.784±.003 3.085±.134 9.499±.184
MLD mld 0.425±.145 0.468±.005 0.656±.003 0.759±.004 3.266±.019 9.698±.094
ReMoDiffuse remodiffuse 0.125±.142 0.493±.004 0.676±.003 0.775±.003 3.047±.007 9.211±.129
LMM* lmm 0.040±.002 0.525±.002 0.719±.002 0.811±.002 2.943±.012 9.814±.076
MotionLab* motionlab 0.167±- 0.810±- 2.912±- 9.593±-
MotionLCM* motionlcm 0.304±.012 0.502±.003 0.698±.002 0.798±.002 3.012±.007 9.607±.006
MotionCLR* motionclr 0.269±.001 0.544±.001 0.732±.001 0.831±.002 2.806±.014
MotionGPT* motiongpt_nips 0.232±.008 0.492±.003 0.681±.003 0.778±.002 3.096±.008 9.528±.071
BiPO* bipo 0.030±.002 0.523±.003 0.714±.002 0.809±.002 2.880±.009 9.556±.076
StableMoFusion* stable_mofusion 0.098±.003 0.553±.003 0.748±.002 0.841±.002 9.748±.092
MoGenTS* mogents 0.033±.001 0.529±.003 0.719±.002 0.812±.002 2.867±.006 9.570±.077
LAMP* lamp 0.032±.002 0.557±.003 0.751±.002 0.843±.001 2.759±.007 9.571±.069
@@
MotionGPT-2* motiongpt2 Llama3-8B 0.191±.004 0.496±.002 0.691±.003 0.782±.004 3.080±.013 9.860±.026
MotionLLM* motionagent GPT4+Gemma-2B 0.230±.009 0.515±.004 0.801±.004 2.967±.020 9.908±.102
Wang et al.largemotionmodel Llama2-13B 0.166±- 0.519±- 0.803±- 2.964±-
ScaMo* scamo codesize 512-3B 0.617±- 0.443±- 0.627±- 0.734±- 3.340±- 9.217±-
AvatarGPT* avatargpt Llama-13B 0.567±- 0.389±- 0.539±- 0.623±- 9.489±-
MotionGPT* motiongpt_aaai Llama-13B 0.567±- 3.775±- 9.006±-
MotionGPT motiongpt_aaai Phi3-3.8B 0.501±.005 0.396±.002 0.575±.005 0.673±.004 3.724±.012 9.475±.110
@@ A.3 More Details of Datasets
IDEA400 represents a high-quality subset of Motion-X motionx , consisting of a large-scale whole-body motion dataset composed of 12.5K clips and 2.6M frames. This dataset encompasses a diverse array of gestures and detailed pose descriptions. Unlike existing works plan_posture , which meticulously select a test set with text descriptions similar to those found in HumanML3D, we adopt a different approach by randomly sampling 10% of the clips without any specific selection criteria. We argue that this methodology more closely resembles an out-of-distribution (OOD) scenario, wherein the test set features a distribution distinctly different from that of the HumanML3D training set.
HumanML3D humanml3d constitutes the largest dataset available, providing text descriptions alongside body-only motions. It includes a total of 14,616 motion clips and 44,970 text descriptions, with 5,371 unique words present across these descriptions. The dataset is partitioned into a training set (80%), a validation set (5%), and a test set (15%).
Regarding pose representation, we adhere to the same specifications as outlined in HumanML3D humanml3d . It is important to note that the number of joints ($J$) is set at 22 for both HumanML3D and IDEA400. We process the motion features for IDEA400 using the same settings as HumanML3D, resulting in a total dimension of 263. These features include root angular velocity, root linear velocities, root height, local joint positions (velocities), and 6D rotations.
@@ A.4 More Experimental Results
We analyze the latency of the VimoRAG framework, as depicted in Table 6, where we report the average time taken per instance during the inference phase. The entire pipeline is divided into three stages: the retrieval phase (conducted by the text-to-video model), the generation phase (performed by the LLM), and the decoding phase (executed by the VQ-VAE). The results indicate that the generation phase exhibits the highest latency, accounting for 92% of the total processing time. This is primarily due to the need for the LLM to load a substantial number of parameters (3.8 billion in VimoRAG). In comparison, the retrieval and decoding phases account for only 6% and 2% of the total time, respectively. This finding highlights that, despite the generation of high-fidelity motion leveraging the world knowledge of the LLM, it remains the latency bottleneck in the overall generation framework. Consequently, this drives us to explore more efficient generative models in our future work. To illustrate the impact of different LLM sizes on latency, we also test the latency of the MotionGPT-13B model motiongpt_aaai on the same computational hardware. As shown in Table 6, the total time for the MotionGPT-13B is nearly twice that of VimoRAG.
@@
In order to investigate the impact of key hyper-parameters on the performance of McDPO training, we conduct comparative experiments focusing on the LoRA fine-tuning parameters rankrank and $\alpha$. As illustrated in Table 1, varying rankrank and $\alpha$ significantly influence the outcomes, a phenomenon that aligns with previous findings in the works motiongpt_aaai ; motiongpt2 . We also observe that when maintaining a constant ratio between rankrank and $\alpha$, both the FID and MultiModal Dist metrics gradually decrease as rankrank and $\alpha$ increase. The increase in rankrank and $\alpha$ corresponds to a greater number of parameters available for optimization within the model, thereby enhancing its capability to fit the dataset more effectively. However, we note a corresponding decline in diversity, indicating that a larger parameter scale adversely affects the model’s ability to generate diverse outputs.
@@ Appendix C Details of Human-centric Video Database
UCF101 ucf-101 13,320 Action Recognition
NTU RGB+D nturgb 114,480 Action Recognition
ASLAN aslan 3,697 Action Similarity Labeling
HMDB51 hmdb51 6,849 Human Motion Recognition
Kinetics-400 kinetics-400 306,245 Human Action Classification
PennAction penn_action 2,326 Action Classification, Action Detection
MotionX motionx 32,500 Human Mesh Recovery, Human Mesh Generation
@@ Appendix D More Implementation Details of VimoRAG
The temporal encoder consists of 4 transformer layers, each featuring 12 attention heads and a width of 768. We utilize learnable position embeddings with a context length of 77 in the action encoder. For the keypoints encoder, a projection layer is placed atop MotionBERT motionbert , with the input and output channels of the projection layer set at 8704 and 768, respectively. To derive the final representation of the keypoints, we implement mean pooling over all the encoded frames, in accordance with existing works clip4clip ; internvideo . The similarity integrator is implemented as a linear transformation, with an input channel of 768 and an output channel of 2. The $\mathcal{L}_{a2p}$ is defined as follows:
@@
In our framework, we employ the VFM InternVideo2 internvideo2 as the video encoder and CLIP-large clip as the image encoder following Maaz et al. videogpt+ . Specifically, we adopt the InternVideo2-Stage2_1B-224p-f4 variant of InternVideo2 and the CLIP-ViT-L/14-336 model. The visual projector is a two-layer MLP with a hidden size of 1024, where a GELU gelu activation function is applied after the first linear layer. For the LoRA parameters, we configure the rank to 128 and set alpha to 256. For the model details and training configurations of VQ-VAE, we adopt the same settings as those used in existing works t2m-gpt ; motiongpt_aaai . Specifically, the codebook size is set to 512x512. The temporal downsampling rate is set to 3 in the encoder of VQ-VAE.
@@
Firstly, we use the $\pi_{ref}$ model obtained from Stage 1 to sample $k$ times (where $k=3$) on a random 25% subset of the training set. The reason for extracting this subset is to reduce inference costs (which is a similar approach used by Zhang et al. llava-hound ). For each input sample, we obtain $k$ different outputs as candidate samples. To identify positive and negative samples from these multiple outputs, we utilize Equation  4 to calculate the reward scores and select the cases with the highest and lowest scores as the positive and negative samples, respectively. We configure $w_{\ell}=0.9$ and $w_{d}=0.1$ in Equation 4 in our experiments.
Appendix E Limitations
Appendix F Impact Statement
Appendix G Future work
In future work, we aim to advance along two main directions. First, we will explore which types of LLMs are most suitable as the backbone of VimoRAG, and how to define appropriate metrics for automatically selecting them. While numerous LLMs are available today, our focus in this paper is not on identifying the best-performing LLM, but rather on addressing the core challenges within the RAG system itself. Second, building upon the success of video-based RAG, we plan to incorporate video, 3D data, and potentially even image data as priors, with the goal of developing a unified RAG framework and investigating whether this multimodal integration can lead to further performance gains.

### v2

@@ 4.1 Datasets, Metrics and Baselines
We conduct extensive experiments on two widely used large-scale datasets following the existing works [16]. The first is the IDEA400 dataset, a high-quality whole-body motion dataset composed of 12.5K clips and 2.6M frames in MotionX [11], which is utilized to assess OOD performance. The other dataset, HumanML3D [5], comprising 14,616 motion clips and 44,970 text descriptions, is utilized to evaluate in-domain performance. For evaluation metrics, we adopt several widely recognized measures: Frechet Inception Distance (FID), R-recall, MultiModal Distance (MM Dist), Diversity. More details can be found in Appendix A due to the space limit.
@@ 4.3 Main Results
Tables 1 and 2 present a quantitative comparison between VimoRAG and SoTA techniques. For a fair comparison, each experiment is conducted ten times, and we report the results with a 95% confidence interval. As illustrated in Table 1, VimoRAG achieves the best FID score, indicating its strong generalization capability in generating high-fidelity motions in OOD scenarios. According to Table 2, VimoRAG outperforms MotionGPT [3] by a large margin across all metrics when using the same backbone. It achieves the best FID score and competitive performance on other metrics among motion LLMs based on the Phi-3 3.8B backbone. As shown in Figure 7, performance improves steadily with larger retrieval sets, highlighting VimoRAG’s potential to enhance motion LLMs.
@@ 4.4 Ablation Study
Impact of Motion Video Retriever.
To study the the influence of the video retriever, we replace the Gemini-MVR with two other retrievers: one being the fine-tuned InternVideo [39] model, and the other being random retrieval from the HcVD. The results in Table LABEL:tab:ablation show that random retrieval leads to a notable increase in the FID score, indicating the importance of accurate video priors in generating high-fidelity results. Moreover, replacing the retriever with InternVideo also leads to a rise in FID score, further confirming the effectiveness of Gemini-MVR. It is important to note that the performance of the retrieval model in RAG is influenced by the generation module, hence we also conduct text-to-video retrieval experiments. As shown in Table LABEL:tab:text-to-video, Gemini-MVR achieves an 8.8% increase in R@1 for the human-centric video set (pool size is 1990), and a 16.6% increase in R@1 for the single human-centric video set. These results further validate the effectiveness of Gemini-MVR.
Impact of McDPO.
@@ 4.5 In-depth Discussion
5 Discussion
Limitations.
Impact Statement.
6 Conclusion and Future Work
@@
In future work, we aim to advance along two directions. First, we will explore which types of LLMs are most suitable as the backbone of VimoRAG, and how to define appropriate metrics for automatically selecting them. While numerous LLMs are available today, our focus in this paper is not on identifying the best-performing LLM, but rather on addressing the core challenges within the RAG system. Second, building upon the success of video-based RAG, we plan to incorporate video, 3D data, and potentially even image data as priors to develop a unified RAG framework and investigating whether this multimodal integration can lead to further performance gains.
Acknowledgments
This work was supported in part by Shenzhen Science and Technology Program (JCYJ20241202123503005). We also acknowledge supports from Guangdong Basic and Applied Basic Research Foundation 2025A1515012281, Nanjing Municipal Science and Technology Bureau 202401035 and University of Macau MYRG-GRG2024-00077-FST-UMDF.
@@ A.1 Experimental Settings
MoMask [13] 0.048±.004 0.519±.005 0.715±.005 0.809±.004 2.955±.011 9.632±.094
T2M-GPT [1] 0.112±.006 0.489±.006 0.679±.006 0.774±.004 3.125±.015 9.691±.062
MDM [20] 0.454±.012 0.419±.004 0.606±.004 0.712±.004 3.636±.015 9.449±.136
MotionDiffuse [2] 0.672±.025 0.492±.004 0.685±.003 0.784±.003 3.085±.134 9.499±.184
MLD [22] 0.425±.145 0.468±.005 0.656±.003 0.759±.004 3.266±.019 9.698±.094
ReMoDiffuse [6] 0.125±.142 0.493±.004 0.676±.003 0.775±.003 3.047±.007 9.211±.129
LMM* [27] 0.040±.002 0.525±.002 0.719±.002 0.811±.002 2.943±.012 9.814±.076
MotionLab* [46] 0.167±- 0.810±- 2.912±- 9.593±-
MotionLCM* [47] 0.304±.012 0.502±.003 0.698±.002 0.798±.002 3.012±.007 9.607±.006
MotionCLR* [48] 0.269±.001 0.544±.001 0.732±.001 0.831±.002 2.806±.014
MotionGPT* [4] 0.232±.008 0.492±.003 0.681±.003 0.778±.002 3.096±.008 9.528±.071
BiPO* [49] 0.030±.002 0.523±.003 0.714±.002 0.809±.002 2.880±.009 9.556±.076
StableMoFusion* [50] 0.098±.003 0.553±.003 0.748±.002 0.841±.002 9.748±.092
MoGenTS* [51] 0.033±.001 0.529±.003 0.719±.002 0.812±.002 2.867±.006 9.570±.077
LAMP* [52] 0.032±.002 0.557±.003 0.751±.002 0.843±.001 2.759±.007 9.571±.069
@@
MotionGPT-2* [25] Llama3-8B 0.191±.004 0.496±.002 0.691±.003 0.782±.004 3.080±.013 9.860±.026
MotionLLM* [53] GPT4+Gemma-2B 0.230±.009 0.515±.004 0.801±.004 2.967±.020 9.908±.102
Wang et al.[54] Llama2-13B 0.166±- 0.519±- 0.803±- 2.964±-
ScaMo* [55] codesize 512-3B 0.617±- 0.443±- 0.627±- 0.734±- 3.340±- 9.217±-
AvatarGPT* [56] Llama-13B 0.567±- 0.389±- 0.539±- 0.623±- 9.489±-
MotionGPT* [3] Llama-13B 0.567±- 3.775±- 9.006±-
MotionGPT [3] Phi3-3.8B 0.501±.005 0.396±.002 0.575±.005 0.673±.004 3.724±.012 9.475±.110
@@ A.3 More Details of Datasets
IDEA400 represents a high-quality subset of Motion-X [11], consisting of a large-scale whole-body motion dataset composed of 12.5K clips and 2.6M frames. This dataset encompasses a diverse array of gestures and detailed pose descriptions. Unlike existing works [16], which meticulously select a test set with text descriptions similar to those found in HumanML3D, we adopt a different approach by randomly sampling 10% of the clips without any specific selection criteria. We argue that this methodology more closely resembles an out-of-distribution (OOD) scenario, wherein the test set features a distribution distinctly different from that of the HumanML3D training set.
HumanML3D [5] constitutes the largest dataset available, providing text descriptions alongside body-only motions. It includes a total of 14,616 motion clips and 44,970 text descriptions, with 5,371 unique words present across these descriptions. The dataset is partitioned into a training set (80%), a validation set (5%), and a test set (15%).
Regarding pose representation, we adhere to the same specifications as outlined in HumanML3D [5]. It is important to note that the number of joints ($J$) is set at 22 for both HumanML3D and IDEA400. We process the motion features for IDEA400 using the same settings as HumanML3D, resulting in a total dimension of 263. These features include root angular velocity, root linear velocities, root height, local joint positions (velocities), and 6D rotations.
@@ A.4 More Experimental Results
We analyze the latency of the VimoRAG framework, as depicted in Table 6, where we report the average time taken per instance during the inference phase. The entire pipeline is divided into three stages: the retrieval phase (conducted by the text-to-video model), the generation phase (performed by the LLM), and the decoding phase (executed by the VQ-VAE). The results indicate that the generation phase exhibits the highest latency, accounting for 92% of the total processing time. This is primarily due to the need for the LLM to load a substantial number of parameters (3.8 billion in VimoRAG). In comparison, the retrieval and decoding phases account for only 6% and 2% of the total time, respectively. This finding highlights that, despite the generation of high-fidelity motion leveraging the world knowledge of the LLM, it remains the latency bottleneck in the overall generation framework. Consequently, this drives us to explore more efficient generative models in our future work. To illustrate the impact of different LLM sizes on latency, we also test the latency of the MotionGPT-13B model [3] on the same computational hardware. As shown in Table 6, the total time for the MotionGPT-13B is nearly twice that of VimoRAG.
@@
In order to investigate the impact of key hyper-parameters on the performance of McDPO training, we conduct comparative experiments focusing on the LoRA fine-tuning parameters rankrank and $\alpha$. As illustrated in Table 1, varying rankrank and $\alpha$ significantly influence the outcomes, a phenomenon that aligns with previous findings in the works [3, 25]. We also observe that when maintaining a constant ratio between rankrank and $\alpha$, both the FID and MultiModal Dist metrics gradually decrease as rankrank and $\alpha$ increase. The increase in rankrank and $\alpha$ corresponds to a greater number of parameters available for optimization within the model, thereby enhancing its capability to fit the dataset more effectively. However, we note a corresponding decline in diversity, indicating that a larger parameter scale adversely affects the model’s ability to generate diverse outputs.
@@ Appendix C Details of Human-centric Video Database
UCF101 [32] 13,320 Action Recognition
NTU RGB+D [33] 114,480 Action Recognition
ASLAN [28] 3,697 Action Similarity Labeling
HMDB51 [29] 6,849 Human Motion Recognition
Kinetics-400 [30] 306,245 Human Action Classification
PennAction [31] 2,326 Action Classification, Action Detection
MotionX [11] 32,500 Human Mesh Recovery, Human Mesh Generation
@@ Appendix D More Implementation Details of VimoRAG
The temporal encoder consists of 4 transformer layers, each featuring 12 attention heads and a width of 768. We utilize learnable position embeddings with a context length of 77 in the action encoder. For the keypoints encoder, a projection layer is placed atop MotionBERT [38], with the input and output channels of the projection layer set at 8704 and 768, respectively. To derive the final representation of the keypoints, we implement mean pooling over all the encoded frames, in accordance with existing works [40, 39]. The similarity integrator is implemented as a linear transformation, with an input channel of 768 and an output channel of 2. The $\mathcal{L}_{a2p}$ is defined as follows:
@@
In our framework, we employ the VFM InternVideo2 [57] as the video encoder and CLIP-large [37] as the image encoder following Maaz et al. [10]. Specifically, we adopt the InternVideo2-Stage2_1B-224p-f4 variant of InternVideo2 and the CLIP-ViT-L/14-336 model. The visual projector is a two-layer MLP with a hidden size of 1024, where a GELU [58] activation function is applied after the first linear layer. For the LoRA parameters, we configure the rank to 128 and set alpha to 256. For the model details and training configurations of VQ-VAE, we adopt the same settings as those used in existing works [1, 3]. Specifically, the codebook size is set to 512x512. The temporal downsampling rate is set to 3 in the encoder of VQ-VAE.
@@
Firstly, we use the $\pi_{ref}$ model obtained from Stage 1 to sample $k$ times (where $k=3$) on a random 25% subset of the training set. The reason for extracting this subset is to reduce inference costs (which is a similar approach used by Zhang et al. [44]). For each input sample, we obtain $k$ different outputs as candidate samples. To identify positive and negative samples from these multiple outputs, we utilize Equation  4 to calculate the reward scores and select the cases with the highest and lowest scores as the positive and negative samples, respectively. We configure $w_{\ell}=0.9$ and $w_{d}=0.1$ in Equation 4 in our experiments.